# Diffusion Attribution Score: Evaluating Training Data Influence in Diffusion Model

### v2

@@ 3 Methodology
In this section, we propose a novel data attribution method tailored for diffusion models to assess the impact of training samples on the generative process. Revisiting our objective, we aim to evaluate how training samples influence the generation within a diffusion model. Previous work D-TRAK, has applied TRAK designed for discriminative tasks to generative tasks, utilizing the Simple Loss as the output function. In Section 3.1, we critically analyze the shortcomings of employing a loss function as the output function in diffusion models and propose a more suitable output function.
Additionally, the attribution score is contingent upon the task setting and the selected output function. Alterations in the settings, such as the loss function and output function, significantly affect the derivation of the attribution score. Consequently, we introduce a new attribution metric called Diffusion Attribution Score (DAS), specifically developed for generative tasks in diffusion models, detailed in Section 3.2. In section 3.3, we explore methodologies for applying DAS to large-scale diffusion model and discuss strategies to expedite the computation process.
3.1 Rethinking the design of model output function
Reviewing the goal of data attribution within diffusion models, suppose we have a set of training data ${\mathbb{S}}$ and wish to measure the influence of a specific training sample ${\bm{z}}^{(i)}$ on a generated sample ${\bm{z}}^{\text{gen}}$. This task can be approached by addressing the counterfactual question: How would the model output change for ${\bm{z}}^{\text{gen}}$ if we removed ${\bm{z}}^{(i)}$ from ${\mathbb{S}}$ and retrained the model on the subset ${\mathbb{S}}_{\backslash i}$? D-TRAK proposes using $\tau_{\text{D-TRAK}}$ to approximate this difference based on the Simple Loss:
@@ 3.1 Rethinking the design of model output function
$=\mathbb{E}_{{\bm{\epsilon}},t}[||{\bm{\epsilon}}_{\theta}({\bm{x}}^{\text{gen}}_{t},t)-{\bm{\epsilon}}||^{2}]-\mathbb{E}_{{\bm{\epsilon}},t}[||{\bm{\epsilon}}_{\theta_{\backslash i}}({\bm{x}}^{\text{gen}}_{t},t)-{\bm{\epsilon}}||^{2}]\text{,}$ (7)
where ${\bm{\epsilon}}_{\theta_{\backslash i}}$ denotes the noise predictor trained on ${\mathbb{S}}_{\backslash i}$. However, from a distribution perspective, this approach involves an indirect comparison of KL-divergences between the model distributions and the data distribution to indicate the distances between model distributions:
$\tau_{\text{D-TRAK}}({\bm{z}}^{\text{gen}},{\mathbb{S}})^{(i)}\approx D_{\mathrm{KL}}[p_{\theta}({\bm{x}}^{\text{gen}})||q({\bm{x}}^{\text{gen}})]-D_{\mathrm{KL}}[p_{\theta_{\backslash i}}({\bm{x}}^{\text{gen}})\|q({\bm{x}}^{\text{gen}})]\text{.}$ (8)
This indirect comparison may introduce errors when evaluating the differences between $p_{\theta}({\bm{x}}^{\text{gen}})$ and $p_{\theta_{\backslash i}}({\bm{x}}^{\text{gen}})$. For instance, both distributions might approach the data distribution $q({\bm{x}}^{\text{gen}})$ from different directions, yet exhibit similar distances. To enable a more direct comparison, we propose the Diffusion Attribution Score (DAS) to assess the KL-divergence between the predicted distributions:
@@
This approach also mitigates the influence of indirect comparisons. However, the subtraction used here does not constitute any recognized type of distance metric. Furthermore, the matrices ${\bm{\epsilon}}_{\theta}$ and ${\bm{\epsilon}}_{\theta_{\backslash i}}$ which retain the same high dimensionality as the input images ${\bm{x}}^{\text{gen}}$, should not be represented merely by scalar values, whether by average or $L^{2}$ norm, as this leads to a loss of dimensional information. For instance, these matrices might exhibit identical differences across various dimensions, an aspect that scalar representations fail to capture. This dimensional consistency is crucial for understanding the full impact of training data alterations on model outputs.
@@ 3.2 Diffusion Attribution Score
In diffusion models, the generative process involves a series of generative steps, and the model produces outputs at $T$ different timesteps. In this section, we explore methods to approximate Eq. 9 at a specific timestep $t$ without the necessity for retraining the model, which aims to evaluate the influence of training data on the model’s generation.
Linearing Output Function. Computing the output of the retrained model ${\bm{\epsilon}}_{\theta_{\backslash i}}({\bm{x}}^{\text{gen}}_{t},t)$ is computationally expensive. To enhance computational efficiency, we propose linearizing the model output function using its Taylor expansion centered around the final model parameters $\theta^{*}$, simplifying the calculation as follows:
@@
The subscript $t$ indicates the attribution score for the model output at timestep $t$. Consequently, the influence of removing a sample from subset ${\mathbb{S}}_{\backslash i}$ on the model output can now be quantitatively evaluated through the changes in model parameters, which can be measured by the leave-one-out method, thereby significantly reducing the computational overhead.
Estimating the model parameter. Consider using the leave-one-out method, the variation of the diffusion model parameters can be assessed by Newton’s Method (Pregibon, 1981). By removing a training sample ${\bm{z}}^{(i)}$, the counterfactual parameters $\theta^{*}_{\backslash i}$ can be approximated by taking a single Newton step from the optimal parameters $\theta^{*}$:
@@
where ${\bm{\epsilon}}_{\theta^{*}}({\mathbb{S}}_{t},t):=[{\bm{\epsilon}}_{\theta^{*}}({\bm{x}}^{(1)}_{t},t),...,{\bm{\epsilon}}_{\theta^{*}}({\bm{x}}^{(n)}_{t},t)]$ represents the stacked output matrix for the set ${\mathbb{S}}$ at timestep $t$, and ${\bm{R}}_{t}:=\text{diag}[{\bm{\epsilon}}_{\theta}({\bm{x}}^{(i)}_{t},t)-{\bm{\epsilon}}]$ is a diagonal matrix describing the residuals among ${\mathbb{S}}$. A detailed proof of Eq. 13 is provided in Appendix B.
@@
where ${\bm{r}}^{(i)}_{t}$ is the $i$-th element of ${\bm{R}}_{t}$. A detailed proof of Eq 15 in provided in Appendix C.
@@ 3.3 Extend DAS to large-scale diffusion model
Reducing Calculation of Expectations Computing $t$ times the equation specified in Eq. 16 is highly resource-intensive due to the necessity of calculating inverse terms. To simplify this process, we consider using the average gradient $\overline{{\bm{g}}}({\bm{x}})$ and average residual $\overline{{\bm{r}}}$, which allows for a singular computation of Eq. 16 to evaluate the overall influence. However, during averaging, these terms may exhibit varying magnitudes across different timesteps, potentially leading to the loss of significant information. To address this, we normalize the gradients and residuals over the entire generative process before averaging:
@@
Reducing the amount of timesteps The computation of Eq. 17 requires performing back propagation $T$ times to calculate the gradients, which is highly resource-intensive. The need to calculate gradients at multiple timesteps can be effectively reduced by sampling fewer timesteps. This method leverages statistical sampling techniques to estimate gradient behaviors across the generative process while significantly reducing computational overhead.
Reduce Dimension of Gradients The dimension of ${\bm{g}}_{t}({\bm{x}}^{(i)})$ is as large as that of the diffusion model itself, which poses a challenge in calculating the inverse term due to its substantial size. There are strategies to simplify this calculation by reducing the dimensionality of the gradient.
One effective method is to apply the Johnson and Lindenstrauss Projection (Johnson & Lindenstrauss, 1984). This technique involves multiplying the gradient vector ${\bm{g}}_{t}({\bm{x}}^{(i)})\in\mathbb{R}^{p}$ by a random matrix ${\bm{P}}\sim\mathcal{N}(0,1)\in\mathbb{R}^{p\times k}(k\ll p)$, which can preserve inner product with high probability while significantly reducing the dimension of the gradient. This projection method has been validated in previous studies (Malladi et al., 2023; Jacot et al., 2018; Park et al., 2023), demonstrating its efficacy in maintaining the integrity of the gradients while easing computational demands. With the above speed up techniques, we summarize our algorithms in Algorithm 1.
In addition to projection methods, other techniques can be employed to reduce the dimension of gradients in diffusion models. For instance, as noted by Ma et al. (2024), the up-block of the U-Net architecture in diffusion models plays a pivotal role in the generation process. Therefore, we can focus solely on the gradients of the up-block for dimension reduction purposes, optimizing computational efficiency.
Furthermore, when applying large-scale diffusion models to downstream tasks, various strategies have been proposed to fine-tune these models efficiently. One such approach is LoRA (Hu et al., 2022), which involves freezing the pre-trained model weights while utilizing trainable rank decomposition matrices. This significantly reduces the number of trainable parameters required for fine-tuning. Consequently, when attributing the influence of training samples in a fine-tuned dataset, we can selectively use the gradients of these trainable parameters to compute the DAS.
Reducing Candidate Training Sample The necessity to traverse the entire training set when computing the DAS poses a significant challenge. To alleviate this, a practical approach involves conducting a preliminary screening to identify the most influential training samples. Techniques such as CLIP (Radford et al., 2021) or cosine similarity can be effectively employed to locate samples that are similar to the target. By using these methods, we can form a preliminary candidate set and concentrate DAS computations on this subset, rather than on the entire training dataset.
@@ 4 Experiments
In this section, we conducted comparative analyses of our method, Diffusion Attribution Score (DAS), against existing data attribution methods under various experimental settings. The primary metric used for assessing attribution performance is the Linear Datamodeling Score (LDS) (Ilyas et al., 2022). Additionally, we evaluated the effectiveness of the speed-up techniques discussed in Section 3.3. Our findings indicate that DAS significantly outperforms other methods in terms of attribution performance, confirming its capability to accurately identify influential training samples.
@@ 4.1 Datasets and Models
This subsection provides an overview of datasets and diffusion models utilized in our experiments. A detailed description of each dataset and model configuration is available in the Appendix F.1.
CIFAR10 (32$\times$32). The CIFAR-10 dataset (Krizhevsky, 2009) consists of 50,000 training samples distributed across 10 classes. We conducted experiments using the entire CIFAR-10 dataset to evaluate the DAS. For computational efficiency and to facilitate ablation studies, we also created a subset called CIFAR-2 from CIFAR-10, which includes 5,000 samples randomly selected from the ”automobile” and ”horse” categories. We employed a DDPM model as described by Ho et al. (2020), which includes 35.7M parameters. Image generation during inference was performed using a 50-step Denoising Diffusion Implicit Models (DDIM) solver (Song et al., 2021a).
@@
ArtBench (256$\times$256). ArtBench (Liao et al., 2022) is specifically curated for generating artwork and features 60,000 images across 10 unique artistic styles. For our studies, we derived ArtBench-2, comprising 5,000 samples from ”post-impressionism” and ”ukiyo-e” styles, and ArtBench-5, including 12,500 samples from five styles including ”post-impressionism”, ”ukiyo-e”, ”romanticism”, ”renaissance” and ”baroque”. We have also created ArtBench-2, which includes 5,000 training samples taken from the 10,000 training in the ”post-impressionism” and ”ukiyo-e” categories. Additionally, we evaluate our method using ArtBench-5, another subset of ArtBench, which is composed of 12,500 training samples drawn from 25,000 training in the ”post-impressionism”, ”ukiyo-e”, ”romanticism”, ”renaissance”, and ”baroque” categories. We fine-tuned a Stable Diffusion model on this dataset using LoRA (Hu et al., 2022) with a rank of 128, involving 25.5M parameters. During inference, new images were generated using a 50-step DDIM solver with a classifier-free guidance scale of 7.5 (Ho & Salimans, 2021).
@@
DAS 33.90$\pm$0.69 43.08$\pm$0.37 20.88$\pm$0.27 30.68$\pm$0.76 24.74$\pm$0.41 33.23$\pm$0.35 13.24$\pm$0.51 18.99$\pm$0.47
@@
4.2 Evaluation Method for Data attribution
Various methods are available for evaluating data attribution techniques, including the leave-one-out influence method (Koh & Liang, 2017; Basu et al., 2021) and Shapley values (Lundberg & Lee, 2017). These methods, however, present significant computational challenges in large-scale settings. To overcome these issues, the Linear Datamodeling Score (LDS) (Ilyas et al., 2022) has been developed as an effective way to assess data attribution methods.
Given a model trained $\theta$ on datset ${\mathbb{S}}$, LDS evaluates the effectiveness of a data attribution method $\tau$ by initially sampling a sub-dataset ${\mathbb{S}}^{\prime}\subset{\mathbb{S}}$ and retraining a model $\theta^{\prime}$ on ${\mathbb{S}}^{\prime}$. The attribution-based output prediction for an interested sample ${\bm{z}}^{\text{test}}$ is then calculated as follows:
where $\rho$ denotes the Spearman rank correlation, ${\mathbb{S}}^{m}$ is the $m$-th subset and $\theta_{m}$ is the model trained on that subset.
In our evaluation, we adopt the model output function setup from D-TRAK (Zheng et al., 2024), setting $f({\bm{z}}^{\text{test}},\theta)$ as the Simple Loss described in Eq. 3 for fair comparison. Although the output functions differ between D-TRAK $f_{\text{D-TRAK}}({\bm{z}},\theta)$ and our DAS $f_{\text{DAS}}({\bm{z}},\theta)$, both methods provide consistent rankings among different subsets. We elaborate on the rationale behind the choice of model output function in LDS evaluations in Appendix E.
For the LDS setting, we set $M=64$ in accordance with D-TRAK’s methodology, with each subset containing half of the entire dataset’s size. For the CIFAR-2, ArtBench-2, and ArtBench-5 datasets, we train three models per subset using different random seeds to ensure robustness. Conversely, for the CIFAR-10 and CelebA datasets, a single model is trained per subset. Each experimental setup involves creating a validation set from the original test set and a generated dataset, each containing 1,000 samples, which serve as ${\bm{z}}^{\text{test}}$ for LDS calculations. Model outputs for validation and generation samples across these datasets are derived using three random seeds. Further details on these configurations are available in the Appendix F.1. Additionally, a comprehensive description of the LDS benchmarks is provided in Appendix F.2.

### v4

@@ 4 Methodology
4.1 Rethinking output function in diffusion models
For the goal of data attribution within diffusion models, we want to measure the influence of a specific training sample ${\bm{z}}^{(i)}$ on a generated sample ${\bm{z}}^{\text{gen}}$. This task can be approached by addressing the counterfactual question: How would ${\bm{z}}^{\text{gen}}$ change if we removed ${\bm{z}}^{(i)}$ from ${\mathbb{S}}$ and retrained the model on the subset ${\mathbb{S}}_{\backslash i}$? This change can be evaluated by the distance between predicted distribution $p_{\theta}({\bm{x}}^{\text{gen}})$ and $p_{\theta_{\backslash i}}({\bm{x}}^{\text{gen}})$ in DDPM, where the subscript denotes the deletion of ${\bm{z}}^{(i)}$.
Reviewing D-TRAK, they proposes using $\tau_{\text{D-TRAK}}$ to approximate this difference on the Simple Loss by setting output function as $f_{\text{D-TRAK}}=\mathcal{L}_{\text{Simple}}$. In details, the difference is computed as:
@@ 4.1 Rethinking output function in diffusion models
$=\mathbb{E}_{{\bm{\epsilon}},t}[||{\bm{\epsilon}}_{\theta}({\bm{x}}^{\text{gen}}_{t},t)-{\bm{\epsilon}}||^{2}]-\mathbb{E}_{{\bm{\epsilon}},t}[||{\bm{\epsilon}}_{\theta_{\backslash i}}({\bm{x}}^{\text{gen}}_{t},t)-{\bm{\epsilon}}||^{2}]\text{.}$ (7)
However, from the distribution perspective, this approach conducts an indirect comparison of KL-divergences between the predicted distributions by involving the data distribution $q({\bm{x}}^{\text{gen}})$:
$\tau_{\text{D-TRAK}}({\bm{z}}^{\text{gen}},{\mathbb{S}})^{(i)}$ $\approx D_{\mathrm{KL}}[p_{\theta}({\bm{x}}^{\text{gen}})\|q({\bm{x}}^{\text{gen}})]-D_{\mathrm{KL}}[p_{\theta_{\backslash i}}({\bm{x}}^{\text{gen}})\|q({\bm{x}}^{\text{gen}})]$
$=D_{\mathrm{KL}}[p_{\theta}({\bm{x}}^{\text{gen}})\|p_{\theta_{\backslash i}}({\bm{x}}^{\text{gen}})]-\int[p_{\theta}({\bm{x}}^{\text{gen}})-p_{\theta_{\backslash i}}({\bm{x}}^{\text{gen}})]\log q({\bm{x}}^{\text{gen}})\text{d}x\text{.}$ (8)
Compared to directly computing the KL-divergences between predicted distributions $p_{\theta}$ and $p_{\theta_{\backslash i}}$, D-TRAK involves a Cross Entropy between the predicted distribution shift $p_{\theta}-p_{\theta_{\backslash i}}$ and data distribution $q$, where the crossing term is generally nonzero unless are identical distributions. This setting involves the influence of the data distribution $q$ in the attribution process and may introduce errors when evaluating the differences. We can consider the irrationality of this setting through a practical example: the predicted distributions might approach the data distribution $q({\bm{x}}^{\text{gen}})$ from different directions while training, yet exhibit similar distances. In this case, D-TRAK captures only minimal loss changes, failing to reflect the true distance between the predicted distributions.
To isolate the effect of the removed data on the model, we propose the Diffusion Attribution Score (DAS), conducting a direct comparison between $p_{\theta}$ and $p_{\theta_{\backslash i}}$ by assessing their KL-divergence:
@@
The effectiveness of $\tau_{\text{Square}}$ lies in manually eliminating the influence of $q$; however, this approach has its limitations since the output of diffusion model is a high dimensional feature. Defining the output function as $\mathcal{L}_{\text{Square}}$ or using average and $L^{2}$ norm, treats the latent as a scalar, thereby neglecting dimensional information. For instance, these matrices might exhibit identical differences across various dimensions, an aspect that scalar representations fail to capture. This dimensional consistency is crucial for understanding the full impact of training data alterations on model outputs.
@@ 4.2 Diffusion Attribution Score
Since we define the output change as KL-divergence, which differs from TRAK, $f_{\text{DAS}}$ cannot be directly applied to TRAK and $\tau_{\text{DAS}}$ needs to be specifically derived on diffusion models. In this section, we explore methods to approximate Eq. 9 at timestep $t$ without retraining the model. The derivation is divided into two main parts: First, we linearize the output function, allowing the difference in the output function to be expressed in terms of difference in model parameters. The second part is that approximating this relationship using Newton’s method. By integrating these two components, we derive the complete formulation of DAS to attribute the output of diffusion model at timestep $t$.
Linearizing Output Function. Computing the output of the retrained model ${\bm{\epsilon}}_{\theta_{\backslash i}}({\bm{x}}^{\text{gen}}_{t},t)$ is computationally expensive. For computational efficiency, we propose linearizing the model output function around the optimal model parameters $\theta^{*}$ at convergence, simplifying the calculation as follows:
@@
The subscript $t$ indicates the attribution for the model output at timestep $t$. Consequently, the influence of removing a sample can be quantitatively evaluated through the changes in model parameters, which can be measured by the Newton’s method, thereby reducing the computational overhead.
Estimating the model parameter. Consider using the leave-one-out method, the variation of the model parameters can be assessed by Newton’s Method (Pregibon, 1981). The counterfactual parameters $\theta^{*}_{\backslash i}$ can be approximated by taking a single Newton step from the optimal parameters $\theta^{*}$:
@@
where ${\bm{\epsilon}}_{\theta^{*}}({\mathbb{S}}_{t},t):=[{\bm{\epsilon}}_{\theta^{*}}({\bm{x}}^{(1)}_{t},t),...,{\bm{\epsilon}}_{\theta^{*}}({\bm{x}}^{(n)}_{t},t)]$ represents the stacked output matrix for the set ${\mathbb{S}}$ at timestep $t$, and ${\bm{R}}_{t}:=\text{diag}[{\bm{\epsilon}}_{\theta}({\bm{x}}^{(i)}_{t},t)-{\bm{\epsilon}}]$ is a diagonal matrix describing the residuals among ${\mathbb{S}}$. A detailed proof of Eq. 13 is provided in Appendix A.
@@
where ${\bm{r}}^{(i)}_{t}$ is the $i$-th element of ${\bm{R}}_{t}$. A detailed proof of Eq 15 in provided in Appendix B.
@@ 4.3 Extend DAS to large-scale diffusion model
Calculating Eq. 16 for large-scale diffusion models poses several challenges. The computation of the inverse term is extremely expensive due to the high dimensionality of the parameters. Additionally, gradients must be calculated for all training samples in ${\mathbb{S}}$, further increasing computational demands. In this subsection, we explore techniques to accelerate the calculation of Eq. 16. These methods can be broadly categorized into two approaches: The first focuses on reducing the gradient computation by minimizing the number of expectations and candidate training samples. The second aims to accelerate the computation of the inverse term by reducing the dimensionality of the gradients.
Reducing Calculation of Expectations. Computing $t$ times the equation specified in Eq. 16 is highly resource-intensive due to the necessity of calculating inverse terms. To simplify, we use the average gradient $\overline{{\bm{g}}}({\bm{x}})$ and average residual $\overline{{\bm{r}}}$ over entire generation, enabling a single computation of Eq.16 to assess overall influence. However, during averaging, these terms may exhibit varying magnitudes across different timesteps, potentially leading to the loss of significant information. To address this, we normalize the gradients and residuals over the entire generation before averaging:
@@
Reducing Dimension of Gradients by Projection. The dimension of ${\bm{g}}_{t}({\bm{x}}^{(i)})$ matches that of the amount of diffusion model’s parameter, posing a challenge in calculating the inverse term due to its substantial size. One effective method of reducing the dimension is to apply the Johnson and Lindenstrauss Projection (Johnson & Lindenstrauss, 1984). It involves multiplying the gradient vector ${\bm{g}}_{t}({\bm{x}}^{(i)})\in\mathbb{R}^{p}$ by a random matrix ${\bm{P}}\sim\mathcal{N}(0,1)\in\mathbb{R}^{p\times k}(k\ll p)$, which can preserve inner product with high probability while significantly reducing the dimension of the gradient. This projection method has been validated in previous studies (Malladi et al., 2023; Zheng et al., 2024), demonstrating its efficacy in maintaining the integrity of the gradients while easing computational demands. We summarize our algorithms in Algorithm 1 with normalization and projection.
Reducing Dimension of Gradients by Model Compression In addition to projection methods, other techniques can be employed to reduce the dimension of gradients in diffusion models. For instance, as noted by Ma et al. (2024), the up-block of the U-Net architecture in diffusion models plays a pivotal role in the generation process. Therefore, we can focus on the up-block gradients for dimension reduction purposes, optimizing computational efficiency. Furthermore, various strategies have been proposed to fine-tune large-scale diffusion models efficiently. One such approach is LoRA (Hu et al., 2022), which involves freezing the pre-trained model weights while utilizing trainable rank decomposition matrices. This significantly reduces the number of trainable parameters required for fine-tuning. Consequently, when attributing the influence of training samples in a fine-tuned dataset, we can compute the DAS with gradients on the trainable parameters.
Reducing the amount of timesteps. Computing Eq. 17 requires performing back propagation $T$ times, making it highly resource-intensive. Sampling fewer timesteps can also approximate the expectation and estimate gradient behavior while significantly lowering computational overhead.
Reducing Candidate Training Sample. The necessity to traverse the entire training set when computing the DAS poses a significant challenge. To alleviate this, a practical approach involves conducting a preliminary screening to identify the most influential training samples. Techniques such as CLIP (Radford et al., 2021) or cosine similarity can be effectively employed to locate samples that are similar to the target. By using these methods, we can form a preliminary candidate set and concentrate DAS computations on this subset, rather than on the entire training dataset.
@@ 5.1 Datasets and Models
In this section, we present a comparative analysis of our method, Diffusion Attribution Score (DAS), against existing data attribution methods across various experimental settings. Our findings demonstrate that DAS significantly outperforms other methods in attribution performance, validating its ability to accurately identify influential training samples. We provide an overview of the datasets and diffusion models used in our experiments, with detailed descriptions available in Appendix E.1.
CIFAR10 (32$\times$32). We conduct experiments on the CIFAR-10 (Krizhevsky, 2009), containing 50,000 training samples across 10 classes. For computational efficiency on ablation studies, CIFAR-2, a subset with 5,000 samples randomly selected from 2 classes is proposed. We use a 35.7M DDPM  (Ho et al., 2020) with a 50-step DDIM solver (Song et al., 2021a) on these settings.
@@
ArtBench (256$\times$256). ArtBench (Liao et al., 2022) is a dataset of 50,000 images across 10 artistic styles. For our studies, we derived two subsets: ArtBench-2, with 5,000 samples from 2 classes and ArtBench-5, with 12,500 samples from five styles. We fine-tuned a Stable Diffusion model on these datasets using LoRA (Hu et al., 2022) with 128 rank and 25.5M parameters. During inference, images are generated using a DDIM solver with a classifier-free guidance (Ho & Salimans, 2021).
@@
DAS 33.90$\pm$0.69 43.08$\pm$0.37 20.88$\pm$0.27 30.68$\pm$0.76 24.74$\pm$0.41 33.23$\pm$0.35 15.24$\pm$0.51 23.69$\pm$0.47
5.2 Evaluation Method for Data attribution
Various methods are available for evaluating data attribution techniques, including the leave-one-out influence method (Koh & Liang, 2017; Basu et al., 2021) and Shapley values (Lundberg & Lee, 2017). In this paper, we use the Linear Datamodeling Score (LDS) (Ilyas et al., 2022) to assess data attribution methods. Given a model trained $\theta$ on dataset ${\mathbb{S}}$, LDS evaluates the effectiveness of a data attribution method $\tau$ by initially sampling a sub-dataset ${\mathbb{S}}^{\prime}\subset{\mathbb{S}}$ and retraining a model $\theta^{\prime}$ on ${\mathbb{S}}^{\prime}$. The attribution-based output prediction for an interested sample ${\bm{z}}^{\text{test}}$ is then calculated as:
where $\rho$ denotes the Spearman correlation and $\theta_{m}$ is the model trained on the $m$-th subset ${\mathbb{S}}^{m}$. In our evaluation, we adopt the output function setup from D-TRAK (Zheng et al., 2024), setting $f({\bm{z}}^{\text{test}},\theta)$ as the Simple Loss described in Eq. 3 for fairness. Although the output functions differ between D-TRAK and DAS, $f_{\text{DAS}}$ can also evaluate the Simple Loss change, where we elaborate on the rationale in Appendix D. Further details about LDS benchmarks are described in Appendix E.2.